How do I solve this system using elimination?

-2x+2y+3z=0
-2x-y+z=-3
2x+3y+3z=5

Answer by Alan3354(69443)

How do I solve this system using elimination?

-2x+2y+3z=0
-2x-y+z=-3
2x+3y+3z=5
-----------
Subtract Eqn 2 from 1.
---
Add Eqn 3 to 1.
----
x is eliminated.